Calculate the fraction of light energy reflected from an air-water interface at normal incidence.
Answer #1

Hi,

In this case we are going to assume that the light comes from the air medium and goes to the water medium and, after that, there is only water (this means that there is only one interface)

The property that can tell us the fraction of light energy reflected at normal incidence is the normal reflection coefficient (R), which can be calculated as:

R = (n2 - n1)/(n2 + n1) ; where n1 is the index of refraction of the medium the light comes from while n2 is the index of refraction of the medium the light is going to.

You can find the value of n2 and n1 in any book, specially in this case as the interface is air-water:

n1 = 1.000 and n2 = 1.333

R = (1.333 - 1.000)/(1.333 + 1.000) = 0.143

So, about a 14 % of the light energy will be reflected, which makes sense since a great deal of light can pass from air to water but some of the incident light is reflected (this is something that can be easily tested).

I hope it helps.